Update for the Week Ending November 18, 2016

Site Remediation Continues

On-going activities:

  • Continued the application of water for dust control;
  • Continued the application of foam and BioSolve for Total Volatile Organic Compounds (TVOC), Particulate Matter (PM-10) [dust], and odor control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
  • Continued pumping from the decontamination pads, excavated areas, sumps and dewatering wells to the water treatment plant (WTP);
  • Continued operation of the WTP and discharge to the sewer;
  • Continued shallow excavation and backfilling on the western areas of the Site;
  • Continued lining and loading trucks and shipping impacted soil to Bayshore Soil Management for treatment and disposal:
    • 131 trucks; 2,699.26 tons shipped last week; and
    • 35,967.39 tons shipped to date.
  • Continued crushing and stockpiling reuse material to be used for backfill in the deep excavation on the eastern area of the Site; and
  • Continued deep excavation within the SOE wall under the temporary containment building (TCB).

Recently Completed:

  • Completed the installation of tiebacks through the support of excavation (SOE) sheet pile wall for external bracing:
    • 1 tieback installed last week; and
    • 76 of 76 tiebacks installed.

Upcoming activities:

  • Complete the testing of the installed tiebacks.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no fifteen-minute Action Limit exceedances for Particulate Matter (PM-10) [dust], Total Volatile Organic Compounds (TVOC), or odors during this reporting period.

There were three instances on Friday evening and Saturday morning, November 18 and 19, where benzene levels exceeded the Action Limit during nonworking hours.  Employees were called back to the Site to repair breaches and tears in the TCB fabric and to apply additional foam to the material stockpiled in the TCB.   Once these items were addressed, benzene and TVOC levels returned to below the Response Limit.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity

Removing hardscape (asphalt and concrete) for reuse material from the shallow excavation area and applying water for dust control.

Excavating impacted soil in the shallow excavation area.

Offloading select fill for backfilling the shallow excavation area.

Installing (drilling) the last tieback through the northern SOE sheet pile wall.

Excavating impacted soil and installing (drilling) the tieback in the TCB.

Holder No. 1 tank uncovered in the south end of the deep excavation in the TCB.

Loading lined trucks with impacted soil from the deep excavation.

Lined trucks waiting to be loaded with impacted soil in the TCB.

Closing the liners of the loaded trucks prior to washing at the decontamination pad.

Washing the lined and loaded trucks at the decontamination pad prior to leaving the Site.